apply the quadratic formula to find the roots of the given function, and then graph the function.
f(x) = x^2 + 4

Answer by venugopalramana(3286)

apply the quadratic formula to find the roots of the given function, and then graph the function.
f(x) = x^2 + 4
D=B^2-4AC=0-16....
NO REAL ROOTS....2 COMPLEX CONJUGATE ROOTS